Find the equation of a line and write it in function notation.
through (3, -8) and perpendicular to 8y= x-16

please help and thank you!!!

Answer by rothauserc(4718)

the perpendicular line has a negative reciprocal slope, therefore
8y = x-16
y = x/8 -2
our perpendicular line has slope -8 (negative reciprocal of 1/8)
so far we have
y = -8x +b
we know the perpendicular line goes through (3, -8)
substitute for x and y
-8 = -8(3) +b
b -24 = -8
b = 16
therefore the equation of the line is
y = -8x +16